mathisboring792:

If a = 3 inches, what is the volume of the paperweight? A. 108 cubic inches B. 81 cubic inches C. 54 cubic inches D. 126 cubic inches

jhonyy9:

what is this ,,a" ?

jhonyy9:

@XioGonz

supie:

I think `A` means `Area` @jhonyy9

supie:

Is there any missing information?

mathisboring792:

no

supie:

@jhonyy9 do you can help ? ty

XioGonz:

Idk..I need more information..

supie:

Is there like a picture or anything you didn't include? @mathisboring792

mathisboring792:

yeah

mathisboring792:

1 attachment
mathisboring792:

there

mathisboring792:

hurry please

unwantedchxd:

a is not area it is apothem and it means the height

jhonyy9:

so there is a cube and a rectangular prisme you need calcule volume of cube with side length ,,a" and the volume of rectangular prism with square base with side length ,,a" and height equal 2a

jhonyy9:

formula for volume of a cube = ?

jhonyy9:

=a^3

jhonyy9:

and volume of a rectangular prisme = ?

jhonyy9:

= area of base time height

mathisboring792:

X^3

jhonyy9:

= a^2 *2a = 2a^3

jhonyy9:

Volume total = a^3 +2a^3 = ?

jhonyy9:

=3a^3 and now you know that a= 3 inches so will be = 3*3^3 = ?

jhonyy9:

3^3 = 3*3*3 = 27 3*27 = ?

mathisboring792:

thanks!

jhonyy9:

= 81 in^3 clearly ?

jhonyy9:

np yw anytime @supie questions ?

supie:

No. ty

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) Great Job!!

OfficialBot:

A=?